OpenStudy (anonymous):

can some one explain how to do this? Solve: | m + 8| = 16 A. {8} B. {-24, 8} C. {8, 16} D. {-24}

OpenStudy (saifoo.khan):

| m + 8| = 16` m+8 = 16 , m + 8 = -16 SOlve both

OpenStudy (anonymous):

what saifoo said

OpenStudy (slaaibak):

what satellite said

OpenStudy (anonymous):

ok i did know what do i do